. A concrete highway is built of slabs 12 m long (20°C). How wide should the
expansion cracks between the slabs be (at 20°C) to prevent buckling if the range of
temperature is -30°C to +50°C?
1
Expert's answer
2019-05-07T09:52:04-0400

Expansion occurs when


"T>20\u00b0C"

The expansion:


"\Delta l=\alpha l (T-T_0)"

"\Delta l=(12\cdot 10^{-6})(12) (50-20)=0.00432 \ m"

The expansion cracks between the slabs should be not less than 4.32 mm.